# Elliptic curves of conductor $2^{m}p$, quadratic twists, and Watkins’s conjecture

### v1

@@ Abstract.
Let $\mathsf{E}/\mathbb{Q}$ be an elliptic curve. By the modularity theorem, it admits a surjection from a modular curve $X_{0}(N)\to\mathsf{E}$, and the minimal degree among such maps is called the modular degree of $\mathsf{E}$. By the Mordell–Weil Theorem, $\mathsf{E}(\mathbb{Q})\simeq\mathbb{Z}^{r}\oplus T$ for some nonnegative integer $r$ and some finite group $T$. Watkins’s Conjecture predicts that $2^{r}$ divides the modular degree, thus suggesting an intriguing link between these geometrically- and algebraically-defined invariants. We offer some new cases of Watkins’s Conjecture, specifically for elliptic curves with additive reduction at $2$, good reduction outside of at most two odd primes, and a rational point of order two.
@@ 1. Introduction
A famous conjecture of M. Watkins [Wat02] predicts that $\mathrm{ord}_{2}(m_{\mathsf{E}})$ is bounded below by the Mordell–Weil rank of $\mathsf{E}(\mathbb{Q})$. Much progress has been made on this problem, especially in the case when $m_{\mathsf{E}}$ is odd (in which case Watkins’s conjecture implies that $\mathsf{E}(\mathbb{Q})$ is finite); see for example [CE09, KK18, KK19, Yaz11].
When $m_{\mathsf{E}}$ is potentially even, much progress has been made in proving Watkins’s Conjecture in quadratic twist families of elliptic curves. For instance, in [EP21], the authors show that if $\mathsf{E}(\mathbb{Q})[2]\neq 0$, then Watkins’s conjecture holds for quadratic twists of $\mathsf{E}$ by square-free integers with sufficiently many prime divisors. In [CP22] the authors establish Watkins’s conjecture for many semi-stable elliptic curves with $\mathsf{E}(\mathbb{Q})[2]\neq 0$ under some additional restrictions on the primes of split and non-split multiplicative reduction. Most recently, J. Caro [Car24] proves that if $\mathsf{E}$ is an elliptic curve with prime-power conductor and $\mathsf{E}(\mathbb{Q})[2]\neq 0$, then Watkins’s Conjecture holds for any quadratic twist of $\mathsf{E}$.
Note that Watkins’s conjecture predicts that $2\mid m_{\mathsf{E}}$ whenever $\mathsf{E}(\mathbb{Q})$ is infinite. As explained in [Car24], the only missing case for this weaker conjecture is the case when $N$ is divisible by at most two odd primes, $\mathsf{E}$ has additive reduction at $2$, and $\mathsf{E}(\mathbb{Q})[2]$ is nontrivial. The results of [Car24] cover the case when $\mathsf{E}$ has additive reduction at $2$ and at the odd primes, but it does not allow for any odd primes of multiplicative reduction.
The goal of this paper is to complement the existing work in this area by establishing Watkins’s conjecture for many elliptic curves with additive reduction at $2$, bad reduction at no more than 2 odd primes, and $\mathsf{E}(\mathbb{Q})[2]\neq 0$. To avoid overlapping with [Car24], we study elliptic curves with multiplicative reduction at one odd prime. We note that the curves we consider often have split multiplicative reduction at an odd prime, so they are not covered by main result of [CP22].
@@
  • We begin with the complete classification, due to W. Ivorra, of elliptic curves of conductor $2^{m}p$ for $m\geq 2$ and $p$ an odd prime. In Theorem 3.1, we prove that many curves with conductor $2^{m}p$ have rank at most 1, including all the curves with $2\leq m\leq 5$. In Corollary 3.4, we show that the rank bound implies Watkins’s Conjecture for these curves, assuming either the Birch and Swinnerton-Dyer Conjecture or the finiteness of Shafarevich–Tate groups.
@@ Acknowledgements
The authors thank Daniel Kriz for answering their questions. Partial support for this research was provided to the first-named author by an AMS-Simons Research Enhancement Grant for Primarily Undergraduate Institution Faculty.
@@ 2. Ivorra’s classification
We sort the remaining curves into convenient families, described below, which are amenable to explict $2$-descent arguments. Each family is introduced in the order of first appearance in the statements [Ivo04, Théorèmes 3–8]. The exact correspondence is given in Table 6 in Appendix B.
@@
For an explicit correspondence between our classification types and the curves listed in [Ivo04], please see Table 6 in Appendix B.
@@ 3. Bounding ranks of Ivorra curves
we have $r=\mathrm{rank}_{\mathbb{Z}}{\mathsf{E}}(\mathbb{Q})=\mathrm{rank}_{\mathbb{Z}}{\mathsf{E}}^{\prime}(\mathbb{Q})$. The general method of $2$-descent (see [ALRP08, Proposition 1.1]) provides a naive upper bound
@@
Here, we give the proof when $T=\mathbf{X}$ to illustrate the technique. We refer the reader to Appendix A for the full proof.
@@
Let $\mathsf{E}$ be an elliptic curve as in the statement of Theorem 3.1. Assuming either the Birch and Swinnerton-Dyer Conjecture or that $\Sha(\mathsf{E}/\mathbb{Q})$ is finite, Watkins’s Conjecture holds for $\mathsf{E}$.
@@
If $\mathrm{rank}_{\mathbb{Z}}\mathsf{E}(\mathbb{Q})=0$ the result is trivial, otherwise we have $\mathrm{rank}_{\mathbb{Z}}\mathsf{E}(\mathbb{Q})=1$. Assuming either BSD or finiteness of $\Sha(\mathsf{E}/\mathbb{Q})$, the analytic rank of $\mathsf{E}/\mathbb{Q}$ is also $1$. By the contrapositive of [CE09, Theorem 1.1] the modular degree of $\mathsf{E}$ is divisible by $2$. Hence Watkins’s Conjecture is satisfied. ∎
@@ 3.2. Observations
VIII $k=2$; $p\equiv 5\pmod{8}$ $\textrm{rank}\leq 1$ by Th 3.1 -
@@ XII $k=2$; $-64\equiv x^{4}\pmod{p}$ see note (a)
XII $k=2$; $-64\not\equiv x^{4}\pmod{p}$ $\textrm{rank}\leq 1$ by Th 3.1 -
@@ XVII $k=1$; $p\equiv 1\pmod{8}$ rank 2 possible 18688b $2^{8}\cdot 73$
XVII $k=1$; $p\equiv 3\pmod{8}$ $\textrm{rank}=0$ by Th 3.1 -
@@ XVII $k=2$ see note (b);
XVIII $k=1$ $\textrm{rank}\leq 1$ by Th 3.1
@@ XIX $k=1$; $p\equiv 1\pmod{8}$ rank 2 possible 24832d $2^{8}\cdot 97$
XIX $k=1$; $p\equiv 7\pmod{8}$ $\textrm{rank}\leq 1$ by Th 3.1 -
@@ XX $k=2$ see note (f) -
4. Watkins’s Conjecture in Quadratic Twist Families
@@
Before proceeding, we also note that the non-triviality of $\mathsf{E}(\mathbb{Q})[2]$ has a useful consequence.
@@ 4.1. Modular form preliminaries
We note that Watkins’s conjecture holds (vacuously) for each of these curves.
@@
Note that Watkins’s conjecture holds for each of these curves, this time for less trivial reasons.
@@ 4.2. Quadratic twists of Ivorra curves
Assume that Goldfeld’s Conjecture is true. Then Watkins’s Conjecture is true for 100% of the quadratic twists $\mathsf{E}^{(d)}$.
@@
Assuming Goldfeld’s Conjecture, 100% of the twists $\mathsf{E}^{(d)}$ have analytic rank $\leq 1$. When the rank is zero, the result is trivial, and when the rank is 1, Watkins’s Conjecture holds by [CE09, Theorem 1.1]. ∎
@@
Since $\mathsf{E}\colon y^{2}=f(x)$ is an Ivorra curve, it has a non-trivial $2$-torsion point $P=(X,0)$ with $X\in\mathbb{Q}$. The quadratic twist is given by $\mathsf{E}^{(d)}\colon dy^{2}=f(x)$, so we also have $P\in\mathsf{E}^{(d)}(\mathbb{Q})[2]\neq 0$. Since $\mathsf{E}^{(d)}$ has additive reduction at $2$ and $q$ and multiplicative reduction at $p$, the result now follows from [DJ19, Lemma 8]. ∎
@@
We now leverage the relationship given by equation (4.2) to prove Watkins’s Conjecture for some quadratic twists of Ivorra curves. We begin with the families of curves covered by Lemma 4.10.
@@
We may use the same technique to deduce Watkins’s Conjecture for twists of other Ivorra curves (such as those appearing in Theorem 3.1) in many cases.
@@
Let $\mathsf{E}/\mathbb{Q}$ be an Ivorra curve and $q\geq 5$ be a prime of good reduction of $\mathsf{E}/\mathbb{Q}$. Then Watkins’s conjecture holds for the quadratic twists $\mathsf{E}^{(\pm q)}/\mathbb{Q}$ in the following cases.
@@
Appendix A Complete $2$-descent proof for Theorem 3.1
@@
(A.1) $r\leq\dim_{\mathbb{F}_{2}}\mathrm{Sel}^{(\phi)}({\mathsf{E}}/\mathbb{Q})+\dim_{\mathbb{F}_{2}}\mathrm{Sel}^{(\phi^{\prime})}({\mathsf{E}}^{\prime}/\mathbb{Q})-2.$
@@ XIV $\delta(\mathcal{O})$ $\mathbb{Q}_{2}$ $\mathbb{Q}_{2}$ $\mathbb{Q}_{2}$ $\mathbb{Q}_{2}$
Appendix B Dictionary to go between Ivorra’s paper and our classification types
@@
Table 6. A dictionary for translating between our classificaiton types and the curves listed in Ivorra [Ivo04].

### v2

# Elliptic curves of conductor $2^{m}p$, quadratic twists, and Watkins’ conjecture

@@ Abstract.
Let $\mathsf{E}/\mathbb{Q}$ be an elliptic curve. By the modularity theorem, it admits a surjection from a modular curve $X_{0}(N)\to\mathsf{E}$, and the minimal degree among such maps is called the modular degree of $\mathsf{E}$. By the Mordell–Weil Theorem, $\mathsf{E}(\mathbb{Q})\simeq\mathbb{Z}^{r}\oplus T$ for some nonnegative integer $r$ and some finite group $T$. Watkins’ Conjecture predicts that $2^{r}$ divides the modular degree, thus suggesting an intriguing link between these geometrically- and algebraically-defined invariants. We offer some new cases of Watkins’ Conjecture, specifically for elliptic curves with additive reduction at $2$, good reduction outside of at most two odd primes, and a rational point of order two.
Résumé. Soit $\mathsf{E}/\mathbb{Q}$ une courbe elliptique. D’après le théorème de modularité, il existe une surjection d’une courbe modulaire $X_{0}(N)\rightarrow\mathsf{E}$, et le degré minimal parmi de telles applications est appelé le degré modulaire de $\mathsf{E}$. D’après le théorème de Mordell–Weil, on a $\mathsf{E}(\mathbb{Q})\simeq\mathbb{Z}^{r}\oplus T$ pour un entier $r\geq 0$ et un groupe fini $T$. La conjecture de Watkins prédit que $2^{r}$ divise le degré modulaire, ce qui suggère un lien intrigant entre ces invariants, qui sont définis géométriquement et algébriquement, respectivement. Nous présentons de nouveaux cas de la conjecture de Watkins, en particulier pour les courbes elliptiques ayant réduction additive en $2$, bonne réduction en dehors d’au plus deux nombres premiers impairs, et un point rationnel d’ordre deux.
@@ 1. Introduction
A famous conjecture of M. Watkins [Wat02] predicts that $\mathrm{ord}_{2}(m_{\mathsf{E}})$ is bounded below by the Mordell–Weil rank of $\mathsf{E}(\mathbb{Q})$. Much progress has been made on this problem, especially in the case when $m_{\mathsf{E}}$ is odd (in which case Watkins’ conjecture implies that $\mathsf{E}(\mathbb{Q})$ is finite); see for example [CE09, KK18, KK19, Yaz11].
When $m_{\mathsf{E}}$ is not assumed to be odd, much progress has been made in proving Watkins’ Conjecture in quadratic twist families of elliptic curves. For instance, in [EP21], the authors show that if $\mathsf{E}(\mathbb{Q})[2]\neq 0$, then Watkins’ conjecture holds for quadratic twists of $\mathsf{E}$ by square-free integers with sufficiently many prime divisors. In [CP22] the authors establish Watkins’ conjecture for many semi-stable elliptic curves with $\mathsf{E}(\mathbb{Q})[2]\neq 0$ under some additional restrictions on the primes of split and non-split multiplicative reduction. Most recently, J. Caro [Car24] proves that if $\mathsf{E}$ is an elliptic curve with prime-power conductor and $\mathsf{E}(\mathbb{Q})[2]\neq 0$, then Watkins’ Conjecture holds for any quadratic twist of $\mathsf{E}$.
Note that Watkins’ conjecture predicts that $2\mid m_{\mathsf{E}}$ whenever $\mathsf{E}(\mathbb{Q})$ is infinite. As explained in [Car24], the only missing case for this weaker conjecture is the case when $N$ is divisible by at most two odd primes, $\mathsf{E}$ has additive reduction at $2$, and $\mathsf{E}(\mathbb{Q})[2]$ is nontrivial; see Appendix A for a discussion of this claim. The results of [Car24] cover the case when $\mathsf{E}$ has additive reduction at $2$ and at the odd primes, but it does not allow for any odd primes of multiplicative reduction.
The goal of this paper is to complement the existing work in this area by establishing Watkins’ conjecture for many elliptic curves with additive reduction at $2$, bad reduction at no more than 2 odd primes, and $\mathsf{E}(\mathbb{Q})[2]\neq 0$. To avoid overlapping with [Car24], we study elliptic curves with multiplicative reduction at one odd prime. We note that the curves we consider often have split multiplicative reduction at an odd prime, so they are not covered by the main result of [CP22].
@@
  • We begin with the complete classification, due to W. Ivorra, of elliptic curves of conductor $2^{m}p$ for $m\geq 2$ and $p$ an odd prime. In Theorem 3.2, we prove that many curves with conductor $2^{m}p$ have rank at most 1, including all the curves with $2\leq m\leq 5$. In Corollary 3.5, we show that the rank bound implies Watkins’ Conjecture for these curves, assuming either the Birch and Swinnerton-Dyer Conjecture or the finiteness of Shafarevich–Tate groups.
@@ Acknowledgements
The authors thank Daniel Kriz for answering their questions, and the referee for helpful suggestions which improved the paper. We also thank Antonio Lei for proofreading our French abstract. Partial support for this research was provided to the first-named author by an AMS–Simons Research Enhancement Grant for Primarily Undergraduate Institution Faculty. DK was supported by an AMS–Simons early career travel grant.
@@ 2. Ivorra’s classification
We sort the remaining curves into convenient families, described below, which are amenable to explict $2$-descent arguments. Each family is introduced in the order of first appearance in the statements [Ivo04, Théorèmes 3–8]. The exact correspondence is given in Table 6 in Appendix C.
@@
For an explicit correspondence between our classification types and the curves listed in [Ivo04], please see Table 6 in Appendix C.
@@ 3. Bounding ranks of Ivorra curves
we have $r=\mathrm{rank}_{\mathbb{Z}}{\mathsf{E}}(\mathbb{Q})=\mathrm{rank}_{\mathbb{Z}}{\mathsf{E}}^{\prime}(\mathbb{Q})$. The general method of $2$-descent provides the following naive upper bound on $r$.
Proposition 3.1.
Keep the notation introduced above. Then
@@
More generally, let $\mathsf{E}/\mathbb{Q}$ be any elliptic curve with a non-trivial point of $2$-torsion and let $a$ (resp. $m$) be number of primes of additive (resp. multi- plicative) bad reduction of $\mathsf{E}/\mathbb{Q}$. Then:
(3.2) $r\leq m+2a-1.$
Proof.
This is [ALRP08, Proposition 1.1]. ∎
@@
Here, we give the proof when $T=\mathbf{X}$ to illustrate the technique. We refer the reader to Appendix B for the full proof.
@@
Let $\mathsf{E}$ be an elliptic curve as in the statement of Theorem 3.2. Assuming either the Birch and Swinnerton-Dyer Conjecture or that $\Sha(\mathsf{E}/\mathbb{Q})$ is finite, Watkins’ Conjecture holds for $\mathsf{E}$.
@@
If $\mathrm{rank}_{\mathbb{Z}}\mathsf{E}(\mathbb{Q})=0$ the result is trivial, otherwise we have $\mathrm{rank}_{\mathbb{Z}}\mathsf{E}(\mathbb{Q})=1$. Assuming either BSD or finiteness of $\Sha(\mathsf{E}/\mathbb{Q})$, the analytic rank of $\mathsf{E}/\mathbb{Q}$ is also $1$. By the contrapositive of [CE09, Theorem 1.1] the modular degree of $\mathsf{E}$ is divisible by $2$. Hence Watkins’ Conjecture is satisfied. ∎
@@ 3.2. Observations
VIII $k=2$; $p\equiv 5\pmod{8}$ $\textrm{rank}\leq 1$ by Th 3.2 -
@@ XII $k=2$; $-64\equiv x^{4}\pmod{p}$ see note (a)
XII $k=2$; $-64\not\equiv x^{4}\pmod{p}$ $\textrm{rank}\leq 1$ by Th 3.2 -
@@ XVII $k=1$; $p\equiv 1\pmod{8}$ rank 2 possible 18688b $2^{8}\cdot 73$
XVII $k=1$; $p\equiv 3\pmod{8}$ $\textrm{rank}=0$ by Th 3.2 -
@@ XVII $k=2$ see note (b);
XVIII $k=1$ $\textrm{rank}\leq 1$ by Th 3.2
@@ XIX $k=1$; $p\equiv 1\pmod{8}$ rank 2 possible 24832d $2^{8}\cdot 97$
XIX $k=1$; $p\equiv 7\pmod{8}$ $\textrm{rank}\leq 1$ by Th 3.2 -
@@ XX $k=2$ see note (f) -
4. Watkins’ Conjecture in Quadratic Twist Families
@@
Before proceeding, we also note that the non-triviality of $\mathsf{E}(\mathbb{Q})[2]$ has a useful consequence. Denote by $a_{q}(\mathsf{E})=q+1-\#\tilde{\mathsf{E}}(\mathbb{F}_{q})$.
@@ 4.1. Modular form preliminaries
We note that Watkins’ conjecture holds (vacuously) for each of these curves.
@@
Note that Watkins’ conjecture holds for each of these curves, this time for less trivial reasons.
@@ 4.2. Quadratic twists of Ivorra curves
Assume that Goldfeld’s Conjecture is true. Then Watkins’ Conjecture is true for 100% of the quadratic twists $\mathsf{E}^{(d)}$.
@@
Assuming Goldfeld’s Conjecture, 100% of the twists $\mathsf{E}^{(d)}$ have analytic rank $\leq 1$. When the rank is zero, the result is trivial, and when the rank is 1, Watkins’ Conjecture holds by [CE09, Theorem 1.1]. ∎
@@
Since $\mathsf{E}\colon y^{2}=f(x)$ is an Ivorra curve, it has a non-trivial $2$-torsion point $P=(X,0)$ with $X\in\mathbb{Q}$. The quadratic twist is given by $\mathsf{E}^{(d)}\colon dy^{2}=f(x)$, so we also have $P\in\mathsf{E}^{(d)}(\mathbb{Q})[2]\neq 0$. Since $\mathsf{E}^{(d)}$ has additive reduction at $2$ and $q$, multiplicative reduction at $p$, and good reduction everywhere else, the result now follows from the bound (3.2). ∎
@@
We now leverage the relationship given by equation (4.2) to prove Watkins’ Conjecture for some quadratic twists of Ivorra curves. We begin with the families of curves covered by Lemma 4.10.
@@
We may use the same technique to deduce Watkins’ Conjecture for twists of other Ivorra curves (such as those appearing in Theorem 3.2) in many cases.
@@
Let $\mathsf{E}/\mathbb{Q}$ be an Ivorra curve and $q\geq 5$ be a prime of good reduction of $\mathsf{E}/\mathbb{Q}$. Then Watkins’ conjecture holds for the quadratic twists $\mathsf{E}^{(\pm q)}/\mathbb{Q}$ in the following cases.
@@
Appendix A Missing Cases of (Weaker) Watkins’ Conjecture
A weaker version of Watkins’ conjecture asserts that
(A.1) $02\nmid m_{\mathsf{E}}\Longrightarrow\operatorname{rk}_{\mathbb{Z}}(\mathsf{E}(\mathbb{Q}))=$ 0 .
In the introduction of [Car24], Caro claims that this conjecture is largely settled, with a few restrictive cases remaining. We now clarify this claim.
Proposition A.1.
Let $\mathsf{E}/\mathbb{Q}$ be an elliptic curve of conductor $N$ and odd modular degree $m_{\mathsf{E}}$. Then the missing cases of (A.1) are when all the following conditions are satisfied
  • $N$ is divisible by at most two odd primes,
  • $\mathsf{E}$ has additive reduction at $2$, and
  • $\mathsf{E}$ has a rational point of order $2$.
Proof.
First, since we are assuming that $2\nmid m_{\mathsf{E}}$, it follows from [CE09, Theorem 1.1(1)] that $N$ is divisible by at most two odd primes.
Next, if $4\nmid N$, then [ARS12, Theorem 2.1] asserts that $m_{\mathsf{E}}$ is equal to the congruence number $\delta_{\mathsf{E}}$ of $\mathsf{E}$. Thus $\delta_{\mathsf{E}}$ is odd, but then the theorem of [KK18] implies that $\mathrm{rank}_{\mathbb{Z}}(\mathsf{E}(\mathbb{Q}))=0$. So the missing case is when $4\mid N$, hence $\mathsf{E}$ has additive reduction at $2$.
These deductions put us in case (a) of [CE09, Theorem 1.1(3)], hence $\mathsf{E}(\mathbb{Q})[2]$ is non-trivial. ∎
Appendix B Complete $2$-descent proof for Theorem 3.2
@@ Appendix A Missing Cases of (Weaker) Watkins’ Conjecture
(B.1) $r\leq\dim_{\mathbb{F}_{2}}\mathrm{Sel}^{(\phi)}({\mathsf{E}}/\mathbb{Q})+\dim_{\mathbb{F}_{2}}\mathrm{Sel}^{(\phi^{\prime})}({\mathsf{E}}^{\prime}/\mathbb{Q})-2.$
@@ XIV $\delta(\mathcal{O})$ $\mathbb{Q}_{2}$ $\mathbb{Q}_{2}$ $\mathbb{Q}_{2}$ $\mathbb{Q}_{2}$
Appendix C Dictionary to go between Ivorra’s paper and our classification types
@@
Table 6. A dictionary for translating between our classification types and the curves listed in Ivorra [Ivo04].